n(n+1)(n+2) for every posi- 7. Use mathematical induction to prove that tive integer n.
Answer #1

boone that CKCK +I)) = n(n+1)(n+2). KE! St use mathematical ineluebon for nal LHS = 7.(1+) =27 T = Lts=fits R.H.S - 1.2.3 =21